# HYDRA: Unifying Multi-modal Generation and Understanding via Representation-Harmonized Tokenization

### 16 Mar 2026 v1

@@ Abstract
Figure 1: Representation schemes in native unified multimodal models. (a) Decoupled Encoder (bagel; janus): It employs a VAE and a representation encoder as dedicated encoders for generation and understanding tasks, respectively. (b) Sequential Encoder (xie2025show): It feeds the output of the VAE directly into the representation encoder in a cascaded manner. (c) Single-representation encoder (ma2025unitok; wu2025harmonizing): It adopts a standalone representation encoder to unify representation learning for both understanding and generation tasks. (d) Our Proposed Representation-Harmonized ViT Design: it also leverages a single ViT backbone, while introducing a bottleneck module to harmonize the feature learning processes of understanding and generation tasks.
@@ 1 Introduction
Unifying visual understanding and generation has emerged as a pivotal frontier in multimodal intelligence (bagel; cao2025hunyuanimage; mogao). Native unified multimodal models (UMMs) (wu2025harmonizing; zhou2024transfusion; xie2025show) advocate for direct decoding within a unified parameter space, demonstrating superior synergy over composite UMMs (seed-x; unilip; blip3-o). However, achieving rational unification is hindered by a fundamental representational divergence, stemming from the fact that understanding and generation constitute inverse tasks with conflicting demands: the former necessitates high-level semantic abstractions, whereas the latter requires compact structural primitives for fine-grained synthesis (clip; vae). This intrinsic conflict forces existing frameworks into disjointed, asymmetric designs, significantly increasing architectural complexity and optimization difficulty.
We attribute this dilemma primarily to the structural limitations of existing image tokenizers, which fail to simultaneously satisfy the three critical criteria illustrated in Fig.  1. First, decoupled paradigms that employ separate encoders for understanding and generation  (bagel; cao2025hunyuanimage) inherently lack Unification of Input Representation, relying on disjoint features that sever the synergy between the two tasks. Second, sequential architectures that stack representation encoders atop VAEs (xie2025show; liu2025tuna) theoretically unify the input but compromise the Coherence of Information Flow due to the significant representation mismatch between the generative VAE latent space and the semantic features required by the representation encoder. Third, while utilizing a single shared representation encoder (unitok; wu2025harmonizing; jiao2025unitoken) attempts to solve this, it often suffers from poor Compatibility of Learning Process, where the conflicting objectives of high-frequency detail preservation and semantic abstraction lead to optimization difficulties. Consequently, current methods face an unavoidable trade-off: they either sacrifice generative fidelity, lose semantic alignment, or struggle to converge on a shared representation.
@@
To this end, we reformulate a ViT-based representation encoder (internvl2.5; internvl) into a progressive learner that transitions from a Generation vision transformer (Gen-ViT), which captures structure-preserving primitives for high-fidelity synthesis, to a Semantic vision transformer (Sem-ViT). To unify these distinct objectives within a single model, we introduce the Generation-Semantic Bottleneck (GSB). The GSB is architected around a novel compress-and-reconstruct operation, creating an information bottleneck that fosters both high-level semantic abstraction and detailed generative fidelity. Specifically, GSB first compresses features into a compact low-dimension space to filter out noise component, then reconstructs them to the original space for subsequent semantic distillation. In this manner, the compact features can both encode structured details while maintaining semantic awareness.
Built upon HYDRA-TOK, we present HYDRA, a unified framework that achieves complete architectural and representational unification. Leveraging the coherent visual representations provided by HYDRA-TOK, visual signals are processed as sequences via a dual-head mechanism, employing an autoregressive head for text prediction and a rectified flow matching head for image generation. Extensive experiments confirm that HYDRA achieves superior performance, highlighting the harmony between understanding and generation during joint training. In terms of multimodal understanding, HYDRA outperforms existing native UMMs by an average margin of approximately 10.0 points across eight benchmarks. Meanwhile, it establishes a new benchmark in visual reconstruction with a remarkable rFID of 0.08, providing a robust foundation that facilitates state-of-the-art generation records: 0.86 on GenEval (ghosh2023geneval), 86.4 on DPG-Bench (hu2024ella), and 0.53 on WISE (niu2025wise). Additionally, we find that joint training consistently outperforms separate training for both generation and understanding, validating the effectiveness of our harmonized representation. Our contributions are summarized as follows:
@@ 2.1 HYDRA-TOK
To ensure robust representation learning across the entire hierarchy, we employ semantic self-distillation on both Gen-ViT and Sem-ViT. We align the intermediate features from distinct depths of the student with a frozen, pre-trained ViT (internvl2.5) via cosine similarity maximization:
@@ 2.2 HYDRA
The unified sequence $\mathbf{U}_{\mathtt{in}}$ is processed by the shared LLM backbone. Following standard UMM practices (xie2025show; bagel), we apply a causal attention mask on language tokens and a bidirectional attention mask on visual tokens within the LLM decoder layers. To address the Compatibility of Learning Process between discrete text prediction and continuous image synthesis, we branch the output of the final transformer layer into two specialized heads. The Language Head (autoregressive) functions as a standard linear classifier, predicting the probability distribution over the text vocabulary based on the unified context: $\mathcal{L}_{\mathtt{NTP}}=-\sum\log P(\mathbf{y}_{i}|\mathbf{y}_{<i},\mathbf{U}_{\mathtt{in}})$. Simultaneously, the Vision Head (diffusion) serves as a continuous regression head modulated by time-step embeddings via AdaLN-Zero (dit). This head operates explicitly on the hidden states corresponding to visual tokens ($\mathbf{H}_{\mathtt{LLM}}^{\mathtt{vis}}$) to predict the flow velocity:
@@ 2.3 Training Recipe
To cultivate the harmonized nature of HYDRA, we implement a three-stage progressive training strategy. We initialize HYDRA-TOK via reconstruction on large-scale corpora combined with semantic self-distillation against a teacher ViT (internvl2.5). Subsequent stages incrementally refine the compatibility between understanding and generation (see Appendix D for full data details).
@@
Table 1: Reconstruction quality on the ImageNet-1K (256 $\times$ 256) and MS-COCO 2017 validation sets. , following the evaluation protocol in (yue2025uniflow). indicates that the model is trained strictly on the ImageNet-1.2M dataset (imagenet)
Method ImageNet-1K MS-COCO 2017 PSNR $\uparrow$ SSIM $\uparrow$ rFID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ rFID $\downarrow$ Generative Only Tokenizer SD-VAE (sd) 23.54 0.68 1.22 26.62 0.77 4.26 SD-VAE XL (sdxl) 27.37 0.78 0.67 27.08 0.80 3.93 SD-VAE 3 (esser2024scaling) 31.29 0.87 0.20 31.18 0.89 1.67 FLUX-VAE (flux) 32.74 0.92 0.18 32.32 0.93 1.35 VA-VAE (vavae) 27.96 0.79 0.28 27.50 0.81 2.71 Wan2.2 (wan2.2) 31.25 0.88 0.75 31.10 0.89 3.28 FlowMo-Hi (flowmo) 26.93 0.79 0.56 Unified Tokenizer Tokenflow (qu2025tokenflow) 21.41 0.69 1.37 UniTok (unitok) 27.28 0.77 0.41 UniLIP (unilip) 22.99 0.747 0.79 EMU2 (emu2) 13.49 0.42 3.27 BLIP3-o (blip3-o) 14.71 0.58 3.18 QLIP (qlip) 23.16 0.63 3.21 UAE (uae) 29.65 0.88 0.19 29.47 0.88 2.04 UniFlow (yue2025uniflow) 33.23 0.96 0.26 32.48 0.95 1.88 RAE (rae) 18.05 0.50 2.04 HYDRA-TOK 37.09 0.97 0.12 35.63 0.98 1.05 HYDRA-TOK 36.39 0.96 0.08 35.94 0.96 0.81
@@
Models # Params # Data MME-S AI2D OCRBench MMB RWQA SEED MMMU MMStar Avg. Understanding-only Models Qwen2.5-VL (bai2023qwen) 3B 1B+ 2199.9 81.4 82.8 77.6 65.5 74.0 51.2 56.3 70.9 Qwen2.5-VL (bai2023qwen) 7B 1B+ 2312.0 84.3 88.8 82.8 68.4 77.0 58.0 64.1 75.7 LLaVA-1.5 (liu2023visual) 7B 0.7M 1510.7 55.5 31.8 62.3 54.8 65.8 35.7 33.1 49.1 LLaVA-OV (li2024llava) 7B 3.2M 1998.1 81.4 62.2 80.8 66.3 76.7 48.8 61.7 68.7 Composite Unified Multimodal Models BLIP3-o (blip3-o) 4B N/A 2161.2 - - - 60.4 73.8 46.6 - - TokenFlow-XL (qu2025tokenflow) 14B N/A 1922.1 - - 68.9 56.6 72.6 43.2 - - SEED-X (seed-x) 17B N/A 1457.0 70.0 - 70.1 49.1 65.6 35.6 - - Smaller Native Unified Multimodal Models Show-o (xie2024show) 1.3B 0.7M 1345.1 - - - - 51.5 27.4 - - JanusFlow (ma2025janusflow) 1.3B 1.4M 1560.1 54.2 - 69.4 41.2 70.5 29.3 40.6 - Janus-Pro (chen2025janus) 1.5B 10M 1875.1 64.5 - 75.5 52.6 68.3 36.3 43.1 - Harmon (wu2025harmonizing) 1.5B 3.2M 1476.2 57.0 11.2 67.1 49.8 67.1 38.9 35.3 47.4 Show-o2 (xie2025show) 1.5B 3.2M 1742.3 69.0 24.5 67.4 56.5 65.6 37.1 43.4 53.2 HYDRA 1.5B 3.2M 1877.9 79.4 50.8 76.8 60.7 74.5 39.1 57.0 63.2 Larger Native Unified Multimodal Models Ming-UniVision (huang2025ming) 16B 20M+ 2023.0 82.8 72.4 78.5 - - 40.3 63.7 - BAGEL (deng2025emerging) 14B 15.2M 2388.0 89.2 73.3 85.0 72.8 78.5 55.3 - - MUSE-VL (muse-vl) 7B 10M - 69.8 - 72.1 - 69.1 39.7 49.6 - Janus-Pro (chen2025janus) 7B 10M 1927.1 71.3 - 79.2 58.0 72.1 41.0 48.3 - VILA-U (wu2024vila) 7B 4M 1845.8 - - 66.6 - 59.0 32.2 - - Show-o2 (xie2025show) 7B 3.2M 1920.5 78.6 32.4 79.3 64.7 69.8 48.9 56.6 62.4 HYDRA 7B 3.2M 2068.6 85.1 57.7 82.4 64.7 75.5 49.4 62.3 68.9
@@
Table 3: Comprehensive image generation results on GenEval (ghosh2023geneval), DPG-Bench (hu2024ella), and WISE (niu2025wise). # Data. indicates the number of image-text pairs used for visual generation. refers to fintuning with GPT-4o distilled synthetic dataset (blip3-o). More Visualization can be founded in Appendix G.2.
@@
Models # Params # Data GenEval DPG-Bench WISE Two Obj. Pos. Over. Glob. Rel. Over. Cult. Space Over. Generation-only Models SD3-Med (esser2024scaling) 2B 1B+ 0.94 0.33 0.74 - - 84.08 - - - FLUX.1 [Dev] (flux) 12B 1B+ 0.93 0.68 0.82 82.10 91.10 84.00 0.48 0.62 0.50 Composite Unified Multimodal Models Blip3-o (blip3-o) 4B 30M - - 0.84 - - 79.36 - - 0.50 TokenFlow-XL (qu2025tokenflow) 14B 60M 0.60 0.16 0.55 87.53 90.30 85.02 - - - SEED-X (seed-x) 17B 158M+ 0.58 0.19 0.49 - - - - - - Smaller Native Unified Multimodal Models Show-o (xie2024show) 1.3B 36M 0.52 0.11 0.53 - - 67.27 - - 0.35 Harmon (wu2025harmonizing) 1.5B 100M 0.86 0.74 0.76 - - - 0.38 0.52 0.41 Show-o2 (xie2025show) 1.5B 66M 0.86 0.46 0.73 87.53 90.30 85.02 0.33 0.53 0.39 HYDRA 1.5B 100M 0.95 0.76 0.86 89.65 91.59 85.51 0.49 0.64 0.49 Larger Native Unified Multimodal Models Ming-UniVision (huang2025ming) 16B 1B+ 0.93 0.92 0.85 - - 82.12 - - - BAGEL (deng2025emerging) 14B 1.6B 0.95 0.78 0.88 88.94 90.82 85.07 0.44 0.68 0.52 MUSE-VL (muse-vl) 7B 24M 0.64 0.25 0.57 - - - - - - Janus-Pro (chen2025janus) 7B 144M 0.89 0.79 0.80 86.90 89.32 84.19 0.30 0.49 0.35 VILA-U (wu2024vila) 7B 15M - - - - - - 0.26 0.37 0.31 Show-o2 (xie2025show) 7B 66M 0.87 0.52 0.76 89.00 91.81 86.14 0.40 0.58 0.44 HYDRA 7B 100M 0.97 0.81 0.86 90.22 91.83 86.40 0.52 0.68 0.53

### 17 Mar 2026 v2

@@ Abstract
Figure 1: Representation schemes in native unified multimodal models. (a) Decoupled Encoder (Deng et al., 2025a; Wu et al., 2025a): It employs a VAE and a representation encoder as dedicated encoders for generation and understanding tasks, respectively. (b) Sequential Encoder (Xie et al., 2025a): It feeds the output of the VAE directly into the representation encoder in a cascaded manner. (c) Single-representation encoder (Ma et al., 2025a; Wu et al., 2025d): It adopts a standalone representation encoder to unify representation learning for both understanding and generation tasks. (d) Our Proposed Representation-Harmonized ViT Design: it also leverages a single ViT backbone, while introducing a bottleneck module to harmonize the feature learning processes of understanding and generation tasks.
@@ 1 Introduction
Unifying visual understanding and generation has emerged as a pivotal frontier in multimodal intelligence (Deng et al., 2025a; Cao et al., 2025; Liao et al., 2025b). Native unified multimodal models (UMMs) (Wu et al., 2025d; Zhou et al., 2024; Xie et al., 2025a) advocate for direct decoding within a unified parameter space, demonstrating superior synergy over composite UMMs (Ge et al., 2024; Tang et al., 2025; Chen et al., 2025a). However, achieving rational unification is hindered by a fundamental representational divergence, stemming from the fact that understanding and generation constitute inverse tasks with conflicting demands: the former necessitates high-level semantic abstractions, whereas the latter requires compact structural primitives for fine-grained synthesis (Radford et al., 2021; Kingma et al., 2019). This intrinsic conflict forces existing frameworks into disjointed, asymmetric designs, significantly increasing architectural complexity and optimization difficulty.
We attribute this dilemma primarily to the structural limitations of existing image tokenizers, which fail to simultaneously satisfy the three critical criteria illustrated in Fig.  1. First, decoupled paradigms that employ separate encoders for understanding and generation  (Deng et al., 2025a; Cao et al., 2025) inherently lack Unification of Input Representation, relying on disjoint features that sever the synergy between the two tasks. Second, sequential architectures that stack representation encoders atop VAEs (Xie et al., 2025a; Liu et al., 2025) theoretically unify the input but compromise the Coherence of Information Flow due to the significant representation mismatch between the generative VAE latent space and the semantic features required by the representation encoder. Third, while utilizing a single shared representation encoder (Ma et al., 2025b; Wu et al., 2025d; Jiao et al., 2025) attempts to solve this, it often suffers from poor Compatibility of Learning Process, where the conflicting objectives of high-frequency detail preservation and semantic abstraction lead to optimization difficulties. Consequently, current methods face an unavoidable trade-off: they either sacrifice generative fidelity, lose semantic alignment, or struggle to converge on a shared representation.
@@
To this end, we reformulate a ViT-based representation encoder (Chen et al., 2024b, c) into a progressive learner that transitions from a Generation vision transformer (Gen-ViT), which captures structure-preserving primitives for high-fidelity synthesis, to a Semantic vision transformer (Sem-ViT). To unify these distinct objectives within a single model, we introduce the Generation-Semantic Bottleneck (GSB). The GSB is architected around a novel compress-and-reconstruct operation, creating an information bottleneck that fosters both high-level semantic abstraction and detailed generative fidelity. Specifically, GSB first compresses features into a compact low-dimension space to filter out noise component, then reconstructs them to the original space for subsequent semantic distillation. In this manner, the compact features can both encode structured details while maintaining semantic awareness.
Built upon HYDRA-TOK, we present HYDRA, a unified framework that achieves complete architectural and representational unification. Leveraging the coherent visual representations provided by HYDRA-TOK, visual signals are processed as sequences via a dual-head mechanism, employing an autoregressive head for text prediction and a rectified flow matching head for image generation. Extensive experiments confirm that HYDRA achieves superior performance, highlighting the harmony between understanding and generation during joint training. In terms of multimodal understanding, HYDRA outperforms existing native UMMs by an average margin of approximately 10.0 points across eight benchmarks. Meanwhile, it establishes a new benchmark in visual reconstruction with a remarkable rFID of 0.08, providing a robust foundation that facilitates state-of-the-art generation records: 0.86 on GenEval (Ghosh et al., 2023), 86.4 on DPG-Bench (Hu et al., 2024), and 0.53 on WISE (Niu et al., 2025). Additionally, we find that joint training consistently outperforms separate training for both generation and understanding, validating the effectiveness of our harmonized representation. Our contributions are summarized as follows:
@@ 2.1 HYDRA-TOK
To ensure robust representation learning across the entire hierarchy, we employ semantic self-distillation on both Gen-ViT and Sem-ViT. We align the intermediate features from distinct depths of the student with a frozen, pre-trained ViT (Chen et al., 2024b) via cosine similarity maximization:
@@ 2.2 HYDRA
The unified sequence $\mathbf{U}_{\mathtt{in}}$ is processed by the shared LLM backbone. Following standard UMM practices (Xie et al., 2025a; Deng et al., 2025a), we apply a causal attention mask on language tokens and a bidirectional attention mask on visual tokens within the LLM decoder layers. To address the Compatibility of Learning Process between discrete text prediction and continuous image synthesis, we branch the output of the final transformer layer into two specialized heads. The Language Head (autoregressive) functions as a standard linear classifier, predicting the probability distribution over the text vocabulary based on the unified context: $\mathcal{L}_{\mathtt{NTP}}=-\sum\log P(\mathbf{y}_{i}|\mathbf{y}_{<i},\mathbf{U}_{\mathtt{in}})$. Simultaneously, the Vision Head (diffusion) serves as a continuous regression head modulated by time-step embeddings via AdaLN-Zero (Peebles & Xie, 2023). This head operates explicitly on the hidden states corresponding to visual tokens ($\mathbf{H}_{\mathtt{LLM}}^{\mathtt{vis}}$) to predict the flow velocity:
@@ 2.3 Training Recipe
To cultivate the harmonized nature of HYDRA, we implement a three-stage progressive training strategy. We initialize HYDRA-TOK via reconstruction on large-scale corpora combined with semantic self-distillation against a teacher ViT (Chen et al., 2024b). Subsequent stages incrementally refine the compatibility between understanding and generation (see Appendix D for full data details).
@@
Table 1: Reconstruction quality on the ImageNet-1K (256 $\times$ 256) and MS-COCO 2017 validation sets. , following the evaluation protocol in (Yue et al., 2025). indicates that the model is trained strictly on the ImageNet-1.2M dataset (Russakovsky et al., 2014)
Method ImageNet-1K MS-COCO 2017 PSNR $\uparrow$ SSIM $\uparrow$ rFID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ rFID $\downarrow$ Generative Only Tokenizer SD-VAE (Rombach et al., 2022) 23.54 0.68 1.22 26.62 0.77 4.26 SD-VAE XL (Podell et al., 2023) 27.37 0.78 0.67 27.08 0.80 3.93 SD-VAE 3 (Esser et al., 2024) 31.29 0.87 0.20 31.18 0.89 1.67 FLUX-VAE (Labs et al., 2025) 32.74 0.92 0.18 32.32 0.93 1.35 VA-VAE (Yao et al., 2025b) 27.96 0.79 0.28 27.50 0.81 2.71 Wan2.2 (Wan et al., 2025) 31.25 0.88 0.75 31.10 0.89 3.28 FlowMo-Hi (Shaulov et al., 2025) 26.93 0.79 0.56 Unified Tokenizer Tokenflow (Qu et al., 2025) 21.41 0.69 1.37 UniTok (Ma et al., 2025b) 27.28 0.77 0.41 UniLIP (Tang et al., 2025) 22.99 0.747 0.79 EMU2 (Sun et al., 2024) 13.49 0.42 3.27 BLIP3-o (Chen et al., 2025a) 14.71 0.58 3.18 QLIP (Zhao et al., 2025) 23.16 0.63 3.21 UAE (Fan et al., 2025) 29.65 0.88 0.19 29.47 0.88 2.04 UniFlow (Yue et al., 2025) 33.23 0.96 0.26 32.48 0.95 1.88 RAE (Zheng et al., 2025) 18.05 0.50 2.04 HYDRA-TOK 37.09 0.97 0.12 35.63 0.98 1.05 HYDRA-TOK 36.39 0.96 0.08 35.94 0.96 0.81
@@
Models # Params # Data MME-S AI2D OCRBench MMB RWQA SEED MMMU MMStar Avg. Understanding-only Models Qwen2.5-VL (Bai et al., 2023) 3B 1B+ 2199.9 81.4 82.8 77.6 65.5 74.0 51.2 56.3 70.9 Qwen2.5-VL (Bai et al., 2023) 7B 1B+ 2312.0 84.3 88.8 82.8 68.4 77.0 58.0 64.1 75.7 LLaVA-1.5 (Liu et al., 2023b) 7B 0.7M 1510.7 55.5 31.8 62.3 54.8 65.8 35.7 33.1 49.1 LLaVA-OV (Li et al., 2024b) 7B 3.2M 1998.1 81.4 62.2 80.8 66.3 76.7 48.8 61.7 68.7 Composite Unified Multimodal Models BLIP3-o (Chen et al., 2025a) 4B N/A 2161.2 - - - 60.4 73.8 46.6 - - TokenFlow-XL (Qu et al., 2025) 14B N/A 1922.1 - - 68.9 56.6 72.6 43.2 - - SEED-X (Ge et al., 2024) 17B N/A 1457.0 70.0 - 70.1 49.1 65.6 35.6 - - Smaller Native Unified Multimodal Models Show-o (Xie et al., 2024) 1.3B 0.7M 1345.1 - - - - 51.5 27.4 - - JanusFlow (Ma et al., 2025c) 1.3B 1.4M 1560.1 54.2 - 69.4 41.2 70.5 29.3 40.6 - Janus-Pro (Chen et al., 2025b) 1.5B 10M 1875.1 64.5 - 75.5 52.6 68.3 36.3 43.1 - Harmon (Wu et al., 2025d) 1.5B 3.2M 1476.2 57.0 11.2 67.1 49.8 67.1 38.9 35.3 47.4 Show-o2 (Xie et al., 2025a) 1.5B 3.2M 1742.3 69.0 24.5 67.4 56.5 65.6 37.1 43.4 53.2 HYDRA 1.5B 3.2M 1877.9 79.4 50.8 76.8 60.7 74.5 39.1 57.0 63.2 Larger Native Unified Multimodal Models Ming-UniVision (Huang et al., 2025) 16B 20M+ 2023.0 82.8 72.4 78.5 - - 40.3 63.7 - BAGEL (Deng et al., 2025b) 14B 15.2M 2388.0 89.2 73.3 85.0 72.8 78.5 55.3 - - MUSE-VL (Xie et al., 2025b) 7B 10M - 69.8 - 72.1 - 69.1 39.7 49.6 - Janus-Pro (Chen et al., 2025b) 7B 10M 1927.1 71.3 - 79.2 58.0 72.1 41.0 48.3 - VILA-U (Wu et al., 2024) 7B 4M 1845.8 - - 66.6 - 59.0 32.2 - - Show-o2 (Xie et al., 2025a) 7B 3.2M 1920.5 78.6 32.4 79.3 64.7 69.8 48.9 56.6 62.4 HYDRA 7B 3.2M 2068.6 85.1 57.7 82.4 64.7 75.5 49.4 62.3 68.9
@@
Table 3: Comprehensive image generation results on GenEval (Ghosh et al., 2023), DPG-Bench (Hu et al., 2024), and WISE (Niu et al., 2025). # Data. indicates the number of image-text pairs used for visual generation. refers to fintuning with GPT-4o distilled synthetic dataset (Chen et al., 2025a). More Visualization can be founded in Appendix G.2.
@@
Models # Params # Data GenEval DPG-Bench WISE Two Obj. Pos. Over. Glob. Rel. Over. Cult. Space Over. Generation-only Models SD3-Med (Esser et al., 2024) 2B 1B+ 0.94 0.33 0.74 - - 84.08 - - - FLUX.1 [Dev] (Labs et al., 2025) 12B 1B+ 0.93 0.68 0.82 82.10 91.10 84.00 0.48 0.62 0.50 Composite Unified Multimodal Models Blip3-o (Chen et al., 2025a) 4B 30M - - 0.84 - - 79.36 - - 0.50 TokenFlow-XL (Qu et al., 2025) 14B 60M 0.60 0.16 0.55 87.53 90.30 85.02 - - - SEED-X (Ge et al., 2024) 17B 158M+ 0.58 0.19 0.49 - - - - - - Smaller Native Unified Multimodal Models Show-o (Xie et al., 2024) 1.3B 36M 0.52 0.11 0.53 - - 67.27 - - 0.35 Harmon (Wu et al., 2025d) 1.5B 100M 0.86 0.74 0.76 - - - 0.38 0.52 0.41 Show-o2 (Xie et al., 2025a) 1.5B 66M 0.86 0.46 0.73 87.53 90.30 85.02 0.33 0.53 0.39 HYDRA 1.5B 100M 0.95 0.76 0.86 89.65 91.59 85.51 0.49 0.64 0.49 Larger Native Unified Multimodal Models Ming-UniVision (Huang et al., 2025) 16B 1B+ 0.93 0.92 0.85 - - 82.12 - - - BAGEL (Deng et al., 2025b) 14B 1.6B 0.95 0.78 0.88 88.94 90.82 85.07 0.44 0.68 0.52 MUSE-VL (Xie et al., 2025b) 7B 24M 0.64 0.25 0.57 - - - - - - Janus-Pro (Chen et al., 2025b) 7B 144M 0.89 0.79 0.80 86.90 89.32 84.19 0.30 0.49 0.35 VILA-U (Wu et al., 2024) 7B 15M - - - - - - 0.26 0.37 0.31 Show-o2 (Xie et al., 2025a) 7B 66M 0.87 0.52 0.76 89.00 91.81 86.14 0.40 0.58 0.44 HYDRA 7B 100M 0.97 0.81 0.86 90.22 91.83 86.40 0.52 0.68 0.53